OpenStudy (anonymous):

How would I solve z - z complex conjugate / 1 + z * z complex conjugate (Im) if every z = x + iy and y =/= 0

hartnn (hartnn):

you mean you need to simplify \(\large \dfrac{z-z^*}{1+zz^*}\) ?

hartnn (hartnn):

if z=x+iy, what will be z* =... ? you know ?

OpenStudy (anonymous):

I mean like this: